### v1

# The SWE-Bench Illusion: When State-of-the-Art LLMs Remember Instead of Reason

Abstract
As large language models (LLMs) become increasingly capable and widely adopted, benchmarks play a central role in assessing their practical utility. For example, SWE-Bench Verified has emerged as a critical benchmark for evaluating LLMs’ software engineering abilities, particularly their aptitude for resolving real-world GitHub issues. Recent LLMs show impressive performance on SWE-Bench, leading to optimism about their capacity for complex coding tasks. However, current evaluation protocols may overstate these models’ true capabilities. It is crucial to distinguish LLMs’ generalizable problem-solving ability and other learned artifacts. In this work, we introduce a diagnostic task: file path identification from issue descriptions alone, to probe models’ underlying knowledge. We present empirical evidence that performance gains on SWE-Bench-Verified may be partially driven by memorization rather than genuine problem-solving. We show that state-of-the-art models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions—without access to repository structure. This performance is merely up to 53% on tasks from repositories not included in SWE-Bench, pointing to possible data contamination or memorization. These findings raise concerns about the validity of existing results and underscore the need for more robust, contamination-resistant benchmarks to reliably evaluate LLMs’ coding abilities.
1 Introduction
As Large Language Models (LLMs) become increasingly integrated into software development tools and workflows, the need for rigorous evaluation of their coding capabilities has grown significantly Srivastava et al. (2023). Standardized benchmarks play a central role in this evaluation, offering reproducible metrics across diverse tasks Zhuo et al. (2024); Pan et al. (2024); Bowman & Dahl (2021). Among these, SWE-Bench Verified has emerged as a prominent benchmark for assessing LLMs’ ability to resolve real-world GitHub issues Jimenez et al. (2024); OpenAI (2024). Recent advances have shown LLMs, e.g. Tools + Claude 4 Opus, achieving over 70% Pass@1 accuracy on SWE-Bench, suggesting substantial progress Anthropic (2025).
However, the rapid improvement in benchmark performance raises a fundamental question: to what extent do these gains reflect genuine, generalizable problem-solving abilities versus memorized patterns from training data? This concern is particularly acute because training corpora for major LLMs often include the same public code repositories from which benchmarks like SWE-Bench are constructed. For instance, GitHub repositories are prominently featured in training datasets for LLaMA Touvron et al. (2023), PaLM Chowdhery et al. (2022), and Codex Chen et al. (2021), creating potential overlap between training and evaluation data.
To investigate this question systematically, we focus on a critical component of software engineering tasks: bug localization—the ability to identify which files contain bugs based on issue descriptions. This subtask is fundamental to successful issue resolution Hossain et al. (2024), yet should require genuine understanding of both the problem description and codebase structure. If models can accurately identify buggy file locations using only issue descriptions (without any repository context), this would suggest memorization of issue-file associations rather than reasoning ability.
We hypothesize that there are two types of memorization that could compromise benchmark validity: instance-specific memorization, where models recall exact issue-solution pairs from training data, and repository-bias memorization, where models have unequal familiarity with different repositories due to training data imbalances. The first type directly invalidates benchmark results by enabling models to retrieve known solutions rather than demonstrate problem-solving ability. The second type is subtler but equally problematic: while repository knowledge is valuable for LLMs, unequal repository familiarity creates systematic evaluation bias. If benchmarks like SWE-Bench predominantly sample from repositories that models know exceptionally well, they may overestimate general coding proficiency. Success might reflect disproportionate familiarity with specific codebases rather than transferable software engineering capabilities.
To systematically investigate these different memorization patterns, we design controlled experiments where models attempt to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context. Our results reveal concerning patterns: state-of-the-art models like o3-mini OpenAI (2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances, despite lacking the contextual information that should be necessary for this task. Critically, performance drops by about 10% on tasks from the same repositories but not included in the SWE-Bench Verified, suggesting potential data contamination or benchmark-specific memorization rather than general coding proficiency.
To further investigate these patterns, we conduct cross-benchmark analysis comparing model performance on SWE-Bench against tasks from external repositories and different programming languages. Across ten evaluated LLMs, we consistently observe superior performance on SWE-Bench compared to external repositories, with particularly dramatic drops (up to 47 percentage points) when evaluating on repositories using different programming languages like C#. These findings raise serious concerns about the validity of current benchmark evaluations in software engineering and highlight the need for more robust, contamination-resistant evaluation approaches.
Our contributions are:
  1. 1.
    Diagnostic Task Design: We introduce a controlled bug localization task that isolates memorization effects from genuine problem-solving by asking LLMs to identify buggy files without providing repository context.
  2. 2.
    Empirical Evidence: We demonstrate through systematic experiments that current SWE-Bench performance may be inflated by memorization, with models achieving suspiciously high accuracy on context-free bug localization tasks.
  3. 3.
    Cross-Benchmark Analysis: We design a multi-benchmark validation process and provide evidence that performance disparities between SWE-Bench and external repositories suggesting repository-bias memorization rather than generalizable coding abilities, calling attention to evaluation validity concerns in the software engineering domain.
2 Approach
2.1 File Path Identification
To investigate potential memorization in large language models, we isolate a focused subtask: identifying buggy file paths from natural language issue descriptions. In our formulation, models are presented with issue reports from various software engineering benchmarks and asked to predict the file path that would require modification to resolve the described issue. Importantly, models do not receive access to repository structure, code, or metadata, only the issue text and repository name are provided. The model’s prediction is considered correct if it exactly matches one of the file paths in the ground truth (GT) patch. By narrowing down the scope to file identification and remove the repository context provided to the LLMs, we aim to expose memorization artifacts apart from their problem solving abilities. In addition, we recognize that issue descriptions themselves may contain file paths. To control for this, we also report a filtered accuracy metric on instances where such explicit path mentions are not observed, details can be found in Section 3.3.
Figure 1 illustrates the prompt template used for this task. Each prompt includes: (1) the name of the repository and (2) the issue description. The model is instructed to predict a single file path that would need to be modified to resolve the issue. Additionally, the model is asked to provide its reasoning, which enables us to analyze its underlying thought process. We show the prompt template in Figure 1 (an example of SWE-Bench instance is shown in Figure 7). We also show an example demonstrating the evaluation process in Figure 2. Here, there is only one GT path and the predicted path overlaps with the GT paths, and thus is considered correct.
You will be provided with an issue statement explaining a problem to resolve within a codebase. The code base is: scikit-learn/scikit-learn.
¡issue¿
Issue description here
¡/issue¿
DISCUSSION
Discuss here with yourself about how you came up with this response.
RESPONSE
path/to/buggy/file.py
File Path Identification Prompt Template
Figure 1: The prompt template used to test models’ ability to identify buggy files without repository access.
Ground Truth Paths: [sklearn/linear_model/ridge.py] Predicted Paths: [sklearn/linear_model/ridge.py] Is_correct: True Evaluation Demonstration
Figure 2: Example to show the evaluation process.
3 Experiment Setup
Here, we introduce the benchmarks we used for our experiment, the metrics we used to evaluate models’ performance on the path identification task, and the settings we used for the LLMs.
3.1 Benchmarks
We include three benchmarks in our preliminary experiments: SWE-Bench-Verified, SWE-Bench-C#, and RefactorBench, which we introduce below.
3.1.1 SWE-Bench-Verified
SWE-Bench-Verified is a human-validated subset of the original SWE-Bench dataset, consisting of 500 curated samples from 12 open-source Python repositories. Each task presents a GitHub issue description and requires generating code edits to resolve it. In standard evaluation, models receive both the repository snapshot and issue description. However, in our path identification experiments, we provide only the issue description and evaluate whether models can correctly identify the file paths that need modification, which should require understanding the repository structure and codebase.
Figure 3: Benchmarks’ repository distributions.
Benchmarks # of Files Issue Length
SWE-Bench Verified 763.5 451.2
SWE-Bench C# 716.7 586.2
RefactorBench 1149.2 14.6
Table 1: Average number of files and average issue description length (in tokens) per instance for each benchmark.
3.1.2 SWE-Bench-C#
SWE-Bench-C# follows the SWE-Bench construction pipeline and comprises 77 tasks from 11 C#-dominant repositories. Each task has the same format as SWE-Bench, requiring code edits based on a given issue description and repository snapshot. For our experiment, we use a subset of 40 tasks from the full dataset, and show the repository distribution in Figure 3 (a). This benchmark allows us to test whether performance advantages observed on Python-based SWE-Bench extend to different programming languages, helping to distinguish between language-specific memorization and genuine coding capabilities.
3.1.3 RefactorBench
RefactorBench Gautam et al. (2025) is a benchmark of 100 handcrafted, multi-file refactoring tasks from nine diverse open-source Python repositories. Each task presents natural language instructions and requires the model to generate a patch that performs the described refactor of the repository. For our experiment, we use a subset of 39 tasks from the full dataset, and show the repository distribution in Figure 3 (b). Unlike SWE-Bench’s focus on bug fixes, RefactorBench tests code restructuring abilities across different repositories and task types. If models show significantly different performance patterns on RefactorBench compared to SWE-Bench despite both being Python-based, this could indicate task-specific optimization or repository-specific optimization rather than general coding proficiency.
In addition, to ensure fair comparison across benchmarks and verify that performance differences reflect contamination rather than varying task difficulty, we analyze two key factors that could influence path identification complexity. In Table 1, we compute the average number of files with target extensions (.py for Python repositories, .cs for C# repositories) as a proxy for search space complexity, and the average issue description length as a measure of available contextual information. As shown in Table 1, SWE-Bench-Verified and SWE-Bench-C# show similar repository sizes (763.5 vs 716.7 files) and comparable issue lengths (451.2 vs 586.2 tokens), suggesting similar task complexity. On the contrary, RefactorBench’s repositories are larger (1149.2 files), and the issue descriptions are dramatically shorter (14.6 tokens), creating a more challenging scenario with larger search spaces but minimal contextual information.
3.2 Additional experiment settings
To more systematically evaluate our benchmark contamination hypothesis, we design a set of complementary test conditions beyond the SWE-Bench-Verified subset. We consider three extensions: (1) randomly sampled tasks from the SWE-Bench-Full dataset, (2) tasks from repositories not present in SWE-Bench but likely included in model training data, and (3) recently created issues from SWE-Bench repositories, which contains instances that postdate SWE-Bench’s construction. These experimental variations enable us to assess both benchmark-specific and temporal memorization patterns.
3.2.1 Full-SWE-Bench
Apart from SWE-Bench-Verified, we additionally evaluate on 200 randomly sampled instances from the full SWE-Bench Jimenez et al. (2024) dataset that are not included in the Verified subset. Since most existing work focuses exclusively on SWE-Bench-Verified, we hypothesize that if performance drops significantly on these unverified instances, it may indicate that models have been optimized specifically for the curated subset or have encountered similar patterns during training, i.e., instance-specific memorization.
3.2.2 Same-Repo Fresh Tasks
As a complementary experiment, we also collected the most recent issues from SWE-Bench repositories that were created after the original SWE-Bench dataset’s cutoff date. This includes tasks from django/django  [^1] https://github.com/django/django (183 tasks), sympy/sympy  [^2] https://github.com/sympy/sympy (119 tasks), and scikit-learn/scikit-learn  [^3] https://github.com/scikit-learn/scikit-learn(42 tasks). These fresh tasks from the same repositories in SWE-Bench provide a critical test: if models perform significantly worse on recent issues from the same repositories where they excel on SWE-Bench tasks, this would strongly suggest that their success stems from memorizing specific historical issues, instance-specific memorization, rather than repository-bias memorization.
3.2.3 External-Repo Tasks
To further test our contamination hypothesis, we constructed tasks from repositories outside the SWE-Bench dataset but with high likelihood of training data inclusion. We collected the earliest available issues and pull requests from two popular repositories: pandas-dev/pandas  [^4] https://github.com/pandas-dev/pandas (123 tasks) and pytorch/pytorch  [^5] https://github.com/pytorch/pytorch (32 tasks). These repositories are widely used, extensively documented, and have been publicly available for years, making them prime candidates for inclusion in large-scale training datasets. This setting helps us recognize repository-bias memorization. If there is strong performance on these extra-repo tasks, this would counter the possibility that models are memorizing the instances on SWE-Bench Verified. However, if models to perform significantly better on SWE-Bench tasks than on these equally-exposed repository tasks, this would indicate that that models are biased toward tasks sourced from repositories included in SWE-Bench, rather than developing general coding capabilities.
3.3 Metrics
We report accuracy as the primary metric to assess the model’s memorization of the task and its familiarity with the corresponding repository.
Accuracy (Acc.)
The accuracy is computed as the percentage of predictions that exactly match the GT patch out of the total number of samples in that category. An exact match is counted when the model-generated file path is identical to at least one of the reference GT file path(s).
Filtered accuracy (F-Acc.)
To address concerns that models might succeed through superficial pattern matching rather than genuine understanding, we introduce filtered accuracy as a control metric. This measures model performance exclusively on issue instances where problem descriptions do not explicitly contain file paths or import statements, identified through simple heuristics. By removing cases where path-like strings appear in issue descriptions, filtered accuracy isolates performance that depends purely on repository knowledge and problem-solving reasoning rather than surface-level string matching. This metric directly addresses our research question of whether models demonstrate genuine coding understanding or rely on memorized patterns.
[…]
3.4 Models
We evaluated a diverse set of ten state-of-the-art LLMs from both OpenAI and Anthropic to measure their ability on file path identification task. From OpenAI, we included snapshot variants of GPT‑4o (gpt‑4o‑2024‑08, and gpt‑4o‑2024‑05), alongside GPT‑4.1 (gpt-4.1-2025-04-14), the reasoning-focused OpenAI o3 (o3-2025-04-16) and its lightweight counterpart o3‑mini (o3-mini-2025-01-31), as well as o4‑mini (o4-mini-2025-04-16), the compact successor OpenAI (2025a). From Anthropic, our suite comprised the Claude family: Claude 3.5 Sonnet (June 20, 2024), Claude 3.7 Sonnet (February 24, 2025), Claude 4.0 Sonnet, and Claude 4.0 Opus, both officially released on May 22, 2025.
For consistency, OpenAI’s chat models and Claude models were configured with a maximum generation length of 2,048 completion tokens. For the reasoning-centric “o” family from OpenAI (i.e., o3, o3‑mini, o4‑mini), we increased the limit to 4,096 tokens. All experiments used the respective provider APIs with default sampling settings to reflect typical usage.
4 Results
4.1 File Path Identification Accuracy
Figure 4: Comparison of model accuracy on the file path identification task across four benchmarks.
Figure 4 presents the file path identification accuracy across all models and benchmarks, revealing a clear hierarchical pattern of performance. These results provide strong evidence for the two types of memorization, instance-specific and repository-bias, hypothesized in our introduction. We analyze the evidence for each pattern below.
4.1.1 Instance-Specific memorization
The first evidence pattern emerges from performance differences within the SWE-Bench repository ecosystem. Models consistently achieve their highest accuracy on the curated SWE-Bench-Verified subset (60-76%), with a noticeable decline on the broader Full-SWE-Bench set (57-71%) and a further drop on newly collected SWE-Repo tasks (50-68%).
This graduated performance decay suggests that models’ knowledge is not uniform even across the same codebases. The superior performance on the ”Verified” set, which is the most commonly used for evaluation and reporting, points toward instance-specific memorization, where models have been disproportionately exposed to or optimized for these specific, widely-circulated benchmark problems.
4.1.2 Repository-Bias memorization
The second, more dramatic pattern appears when comparing performance between SWE-Bench and external benchmarks. While this performance drop is partly compounded by confounding variables, such as the significantly shorter issue descriptions in RefactorBench (Table 1), the data still points overwhelmingly to repository-level bias.
The most compelling evidence comes from the Outside-Repo Tasks. These tasks were sourced from popular repositories (pandas, PyTorch) that are almost certainly well-represented in training data. If models developed a general capability for popular Python projects, we would expect strong performance here. Instead, accuracy is uniformly lower (below 53%) than those in SWE-Bench Verified.
The stark contrast between high performance on the SWE-Bench repositories and poor performance on these other popular repositories refutes the idea that models are simply good at ”popular projects.” It strongly supports our hypothesis of repository-bias memorization, where models have overfit to the specific architectural patterns and problem distributions of the dozen repositories in the SWE-Bench collection, failing to develop transferable skills.
4.1.3 Model-Specific Contamination Patterns
The results reveal differences in contamination patterns across model families and vendors:
OpenAI chat Models: OpenAI’s chat model variants show inconsistent performance patterns across benchmarks, with no stable ranking order. The performance ranking of these models varies significantly between different benchmark types. The only consistent pattern is GPT-4o-2024-08-06, which maintains mediocre performance across most evaluations.
OpenAI reasoning Models: While o3 demonstrates higher accuracy than o3-mini and o4-mini due to its enhanced capabilities, the performance patterns of o3-mini and o4-mini follow similar trends to the chat models. We cannot identify a consistently stronger model within this family for the path prediction task across different benchmarks.
Anthropic Models: Claude models exhibit more consistent performance patterns, with newer model generations generally showing higher path prediction accuracy. However, Claude Sonnet 4 shows notably lower performance than other Claude family models on RefactorBench and Extra-Repo tasks, which represents an exception to the otherwise consistent generational improvement pattern. Overall, Claude Sonnet 4 and Opus 4 maintain relatively higher performance across most benchmarks compared to other Claude models.
Cross-Vendor Consistency: A key finding from our evaluation of ten different LLMs is the remarkable consistency of these patterns across different model families and vendors. The fact that all models, regardless of vendor, exhibit the same performance hierarchy (SWE-bench Verified >>> Full-SWE-Bench >>> SWE-Repo Tasks >>> SWE-bench C#, RefactorBench, and Outside-Repo Tasks) indicates the performance disparities reflect systematic exposure patterns in training data rather than isolated vendor-specific issues.
4.1.4 Implications for Benchmark Validity
The results provide strong empirical evidence for the two forms of memorization hypothesized in our introduction. The performance degradation within the SWE-Bench ecosystem points to instance-specific memorization, while the more significant performance collapse on external repositories indicates repository-bias memorization. The dual presence of these phenomena suggests that high scores on SWE-Bench are not purely indicative of generalizable coding ability. Instead, they are likely inflated by these confounding factors, reflecting a specialized knowledge of the benchmark’s specific instances and repository patterns rather than transferable software engineering skills.
Dataset Mentioned # (%) Unmentioned # (%) Total #
SWE-bench Verified 135 (27.0%) 365 (73.0%) 500
Full-SWE-Bench 42 (21.0%) 158 (79.0%) 200
SWE-Repo Tasks 100 (29.1%) 244 (70.9%) 344
SWE-bench C# 11 (14.7%) 64 (83.1%) 77
RefactorBench 1 (2.6%) 38 (97.4%) 39
Outside-Repo Tasks 26 (16.5%) 132 (83.5%) 158
Table 2: Dataset splits with corresponding proportions (as percentages) for the instances that contains ground truth path information and those do not.
4.2 File Path Identification Accuracy on Filtered Instances
To more rigorously test our contamination hypothesis, this section examines model performance only on instances where the issue description provides no explicit file paths or import statements. The results, presented in Figure 5, remove superficial string matching as a confounding variable and reveal that the underlying memorization patterns persist.
First, models still achieve their highest accuracy on SWE-Bench Verified instances, followed by Full SWE-Bench and SWE-Repo tasks confirming a pattern of instance-specific memorization that is not dependent on clues from the prompt. Second, the sharp performance drop on external repositories also persists, providing definitive evidence for repository-bias memorization.
The persistence of these patterns in the filtered data is our most critical finding. It shows that the models’ ability to identify correct files is not due to simple text matching but is based on memorized knowledge, which includes instance-specific patterns and repository-biased patterns. This strongly suggests that current benchmark scores are inflated by these memorization effects and do not purely reflect a model’s transferable software engineering skills.
Figure 5: Comparison of models’ filtered accuracy on the file path identification task across benchmarks.
5 Related Work
5.1 Coding Benchmarks and Agents
Various coding benchmarks have been proposed to evaluate LLMs and LLM agents’ coding abilities, including SWE-Bench, BigCodeBench, SWE-Gym, EvoEval Jimenez et al. (2024); Zhuo et al. (2024); Xia et al. (2024). Among them, SWE-Bench has emerged as one of the most popular evaluation method for LLM’s coding abilities (Jimenez et al., 2024). The benchmark consists of real GitHub issues and their corresponding patches, providing a seemingly realistic evaluation setting. However, the public nature of GitHub data raises concerns about potential exposure during training. Recent agent systems have shown dramatic improvements on SWE-Bench (Yang et al., 2024; Wang et al., 2024), but the extent to which these improvements reflect genuine problem-solving versus pattern matching remains unclear.
5.2 Benchmark Contamination
The issue of benchmark contamination, where models are inadvertently trained on evaluation data, is a well-documented problem that can lead to inflated performance metrics without genuine understanding Zhou et al. (2023). To specifically combat this in the coding domain, prior work has developed two primary strategies: task mutation and metric-based probing.
Task mutation approaches aim to detect memorization by altering existing benchmark problems. For example, EvoEval Xia et al. (2024) uses LLMs to create semantic-altering transformations of HumanEval Chen et al. (2021) problems, while other methods Chen et al. (2025) use AST-based mutations to measure how performance changes with structural code modifications. Similarly, TaskEval Tambon et al. (2025) measures prompt sensitivity by creating numerous variations of the same task. These methods are effective at revealing overfitting on specific problem formulations.
Metric-based probing approaches, in contrast, analyze the model’s outputs without changing the task itself. These techniques Li (2023); Ramos et al. (2025) include using perplexity analysis or other statistical measures like Negative Log Likelihood (NLL) and n-gram accuracy to infer if a model has likely seen a specific bug or solution before. However, perplexity and NLL approaches cannot be applied to commercial models, for we do not have access to the model’s hidden states. N-gram similarity, while applicable, can be a noisy indicator for complex code patches; any functionally correct solution will inevitably share significant textual overlap with the ground truth, making it difficult to distinguish genuine problem-solving from verbatim memorization.
While existing mutation and metric-based methods are effective for single-function code generation, they primarily target instance-specific memorization and are less suited for complex, multi-file software engineering tasks like those in SWE-Bench. Our work introduces a complementary diagnostic approach that targets a different type of contamination: repository-bias memorization. Through subtask isolation—testing models on prerequisite tasks without necessary context—and systematic cross-benchmark analysis across different repository sets and programming languages, we provide a methodology to detect whether high benchmark performance reflects genuine problem-solving ability or disproportionate familiarity with specific codebases.
6 Conclusion
This work presents evidence that current evaluations of LLM coding capabilities may be compromised by benchmark contamination and data memorization. Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific memorization, evidenced by graduated performance decay within the SWE-Bench ecosystem, and repository-bias memorization, demonstrated by substantial performance drops on external repositories (up to 47 percentage points) despite their popularity and likely training data inclusion. The consistency of these patterns across different model families and vendors indicates systematic exposure patterns rather than isolated issues.
These findings suggest that widely-cited improvements in SWE-Bench performance may partially reflect benchmark-specific optimization rather than genuine advances in coding capabilities. The field urgently needs more robust evaluation frameworks that incorporate temporal controls to prevent training data contamination, cross-repository validation to test generalization beyond familiar codebases, and systematic cross-benchmark analysis to distinguish between memorization and transferable skills. Our work highlights the importance of developing contamination-resistant benchmarks and more sophisticated evaluation methodologies to ensure that reported progress reflects genuine advances in software engineering capabilities rather than dataset-specific artifacts.
Appendix A Appendix
A.1 Full Prompt Template
def get_file_path_prompt(entry):
    file_prompt = ’’’You will be provided with an issue statement explaining a problem to resolve within a codebase. The code base is: {0}.
   <issue>
   {1}
   </issue>
   You must always include one discussion and one response as part of your response.
   Make sure you do not have multiple discussion/response tags.
   Please make sure your output precisely matches the following format.
   DISCUSSION
   Discuss here with yourself about how you came up with this response.
   RESPONSE
   response(s) to the question asked.
   Now answer the following question:
   Given the issue description and project, provide a file-path of the .py file containing the issue relative to the root.’’’.format(entry[’repo’], entry[’problem_statement’])
    model_input = [{
       ’role’: ’user’,
       ’content’: file_prompt
    return model_input
Complete File Path Prompt Implementation
Figure 6: The complete implementation of the file path identification prompt used in our experiments.
RidgeClassifierCV store_cv_values Issue Error: TypeError: __init__() got an unexpected keyword argument ’store_cv_values’ Repro: from sklearn import linear_model import numpy as np x = np.random.randn(100, 30) y = np.random.normal(size=100) model = linear_model.RidgeClassifierCV( alphas=np.arange(0.1, 1000, 0.1), normalize=True, store_cv_values=True ).fit(x, y) Note: store_cv_values is not a valid parameter, yet some attributes depend on it.
Figure 7: Minimal reproduction of RidgeClassifierCV parameter issue
A.2 Additional Results
[Additional tables and figures supporting the main findings]
Model Mentioned Not Mentioned Overall
Single GT Multi GT Overall Single GT Multi GT Overall Single GT Multi GT Overall
claude-3-5-sonnet 88.60 71.43 85.93 51.90 44.90 50.96 61.63 52.86 60.40
claude-3-7-sonnet 92.11 85.71 91.11 60.13 57.14 59.73 68.60 65.71 68.20
claude-sonnet-4 91.23 85.71 90.37 65.51 53.06 63.84 72.33 62.86 71.00
claude-opus-4 92.11 95.24 92.59 67.41 61.22 66.58 73.95 71.43 73.60
gpt-4o (May) 91.23 85.71 90.37 51.27 55.10 51.78 61.86 64.29 62.20
gpt-4o 88.60 80.95 87.41 53.16 51.02 52.88 62.56 60.00 62.20
gpt-4.1 85.09 71.43 82.96 56.65 55.10 56.44 64.19 60.00 63.60
o3-mini 91.23 85.71 90.37 60.13 63.27 60.55 68.37 70.00 68.60
o3 92.98 76.19 90.37 71.84 65.31 70.96 77.44 68.57 76.20
o4-mini 92.98 80.95 91.11 59.49 48.98 58.08 68.37 58.57 67.00
Table 3: Accuracy of model-predicted file paths on SWE-Bench Verified, stratified by whether any GT path is explicitly mentioned in the problem description. Results are further broken down into three categories: Single GT Path (tasks where the GT patch modifies only a single execution path), Multiple GT Paths (tasks involving modifications across multiple execution paths), and Overall (all tasks). Among the task instances, 135 mention at least one GT path, while 365 do not.